# Img2CAD: Reverse Engineering 3D CAD Models from Images through VLM-Assisted Conditional Factorization

### v1

@@
Figure 1. Img2CAD: A Framework for Reverse Engineering 3D CAD Models from Single-View Images. Our method leverages GPT-4V to predict the discrete CAD program structure and then uses a semantic-conditioned transformer to predict the continuous attributes. This approach allows users to easily reconstruct and edit a CAD model from a single-view input image. We use OpenCasCade (Capgemini, [n. d.]) to convert the CAD program back to a 3D mesh.
Img2CAD: Reverse Engineering 3D CAD Models from Images through VLM-Assisted Conditional Factorization
Yang You Stanford UniversityStanfordCA94305USA yangyou@stanford.edu Mikaela Angelina Uy Stanford UniversityStanfordCA94305USA mikacuy@cs.stanford.edu Jiaqi Han Stanford UniversityStanfordCA94305USA jiaqihan@stanford.edu Rahul Thomas Stanford UniversityStanfordCA94305USA rt03mas@stanford.edu Haotong Zhang Peking UniversityBeijingChina 2200012835@stu.pku.edu.cn Suya You The U.S. Army Research LabMarylandUSA suya.you.civ@army.mil  and  Leonidas Guibas Stanford UniversityStanfordCA94305USA guibas@cs.stanford.edu
CAD modeling
journal: TOG
Reverse engineering 3D computer-aided design (CAD) models from images is an important task for many downstream applications including interactive editing, manufacturing, architecture, robotics, etc. The difficulty of the task lies in vast representational disparities between the CAD output and the image input. CAD models are precise, programmatic constructs that involves sequential operations combining discrete command structure with continuous attributes – making it challenging to learn and optimize in an end-to-end fashion. Concurrently, input images introduce inherent challenges such as photo-metric variability and sensor noise, complicating the reverse engineering process. In this work, we introduce a novel approach that conditionally factorizes the task into two sub-problems. First, we leverage large foundation models, particularly GPT-4V, to predict the global discrete base structure with semantic information. Second, we propose TrAssembler that conditioned on the discrete structure with semantics predicts the continuous attribute values. To support the training of our TrAssembler, we further constructed an annotated CAD dataset of common objects from ShapeNet. Putting all together, our approach and data demonstrate significant first steps towards CAD-ifying images in the wild. Our project page: anonymous123342.github.io
@@ 1. Introduction
Secondly, CAD systems, despite their utility in applications requiring detailed editing and manipulation, present significant challenges due to its combination of discrete command types with continuous parameter values. A CAD model is a sequence of discrete geometric operations of different types, e.g. line, arc, circle, extrude add, extrude cut – each characterized by continuous attributes such as length, position, and angle. Learning this dual nature of both discrete operations and their continuous attributes poses a non-trivial challenge, necessitating algorithms to predict both the structural and attribute components accurately. The hybrid discrete-continuous nature of the problem makes image-to-CAD learning and optimization non-trivial. This approach may require an exponential amount of data to learn a combinatorial space of structures, which is difficult to obtain.
@@
To tackle these problems, we propose to conditionally factorize the task into two sub-problems – to first predict the global discrete base structure and then, conditioned on this base structure, to regress the continuous attribute values. For the first part, we leverage the generalizability of large vision foundation models, particularly GPT-4V, to predict the global base structure from an image. We then propose a transformer-based network named TrAssembler, to regress the continuous attributes conditioned on the discreate structure. While the space of possible object structures is vast, even within a single category such as chairs, shared substructures among objects are extremely common. For example, many chairs have support structures consisting of four legs. Moreover, we find that a VLM can generate semantically consistent part labels as program comments for such common substructures. These labels can then be used by the continuous attribution prediction network to effectively establish correspondences that facilitate communication and shared learning across corresponding attributes (e.g., leg length) of different objects, thus enabling the sharing of information across the attribute space. Our conceptual idea is inspired by the mathematical notion of sheaves (Curry, 2014), where fibers (the attribute parameter spaces in our case) over points in a base space (the space of labelled CAD structures in our case) have partial locally consistent structure (through corresponding attributes).
@@
To train TrAssembler, we further introduce a CAD dataset of common objects. This was constructed by leveraging the rich part annotations of PartNet (Mo et al., 2019b) and the general assistance of VLMs – resulting in our CAD-ified subset of ShapeNet (Chang et al., 2015). Our experiments demonstrate the feasibility of our approach for image-to-CAD, which, to the best of our knowledge, is the first of its kind for a general sketch-extrude CAD representation of common objects. We show that our method outperforms baselines and further showcase various applications, such as CAD-ification and shape editing from images in the wild.
@@ 2.1. Reverse Engineering CAD
Reverse engineering in CAD has been a well-studied focus in computer graphics and CAD communities (Bénière et al., 2013; Várady et al., 1997; Benko and Varady, 2004; Benkö et al., 2002; Ma et al., 2023), allowing for the structured representation of shapes for different tasks such as editing and manufacturing. Initial works explored reconstructing CAD through inverse CSG modeling (Du et al., 2018; Sharma et al., 2018; Kania et al., 2020; Ren et al., 2021; Yu et al., 2023) that decomposes shapes into simple volumes combined with boolean operations. Despite being effective for reconstruction, these methods often involve numerous components with a decomposition that is not compact, making them less flexible and harder to be manipulated by users. Recent advancements have been driven by large-scale CAD datasets (ABC (Koch et al., 2019), Fusion 360 (Willis et al., 2021), SketchGraphs (Seff et al., 2020)) that enabled the development of data-driven techniques for B-Rep classification (Jayaraman et al., 2021), segmentation (Jayaraman et al., 2021; Lambourne et al., 2021; Dupont et al., 2022), parametric surface inference (Guo et al., 2022; Liu et al., 2023), CAD reconstruction (Khan et al., 2024; Lambourne et al., 2022) and generation (Xu et al., 2023). Different from these datasets that mostly consist of mechanical parts, one of our contributions is a CAD dataset of common every day objects that contain meaningful semantic parts, allowing for more natural user editing.
In reverse engineering, a prevalent CAD modeling approach employs the sketch-extrude (Willis et al., 2021; Uy et al., 2022) paradigm, where a sketch (Xu et al., 2022; Para et al., 2021) is defined as a series of parametric curves forming a closed loop, resulting in flexible and complex primitives, which is the focus of this work. Some recent works have borrowed ideas from NLP to interpret CAD models as sequences of tokens (Wu et al., 2021; Ganin et al., 2021; Jayaraman et al., 2023; Zhang et al., 2024). Point2Cyl (Uy et al., 2022) takes a different approach and redefines the CAD reverse engineering task as a geometry-aware decomposition problem. It decomposes an input point cloud into extrusion cylinders (sketch-extrude primitives), and subsequent works extended this approach to the unsupervised (Ren et al., 2022) and self-supervised (Li et al., 2023) settings. These works however take in 3D point clouds as input; in contrast, our work tackles a more challenging setting with a single-view image input.
@@ 2.2. Structured Representation of Shapes
Our daily objects are often highly structured in their geometry and part composition. For example, a chair is composed of parts governed by a set of rules and constraints (Ritchie et al., 2023; Uy et al., 2020, 2021), such as having legs of equal lengths placed symmetrically around the seat. Existing works have explored various ways of representing and learning structure from shapes, such as part-based templates (Ganapathi-Subramanian et al., 2018; Kim et al., 2013; Ovsjanikov et al., 2011), symmetries (Mitra et al., 2006; Sung et al., 2015), shape programs (Müller et al., 2023; Jones et al., 2020, 2021, 2023), and grammars (Kalogerakis et al., 2012; Chaudhuri et al., 2011). Notably, GRASS (Li et al., 2017) and StructureNet (Mo et al., 2019a) represent a shape as a graph of part hierarchies learned through a supervised neural network. The goal of this line of work is to learn higher-level semantic structures, i.e., parts of shapes, but does not focus the reconstruction of the local geometric detail of each part. These methods typically assume a coarse and homogeneous representation for each part, commonly as bounding boxes. In contrast, our goal is to address image-to-CAD for daily objects, which requires not only understanding the underlying structure but also reconstructing each part’s geometry.
@@ 2.3. Primitive Fitting
Another relevant area in computer graphics is primitive fitting. One line of work involves decomposing shapes into surface or volumetric primitives. Classical graphics approaches include detecting simple primitives such as planes (Czerniawski et al., 2018; Fang et al., 2018; Monszpart et al., 2015) using methods like RANSAC (Schnabel et al., 2007; Li et al., 2011), region growing (Oesau et al., 2016), or Hough voting (Borrmann et al., 2011; Drost and Ilic, 2015). Data-driven approaches (Li et al., 2019; Sharma et al., 2020; Lê et al., 2021) have also grown in popularity, leveraging models that fit simple primitives such as planes, cylinders, cones, and spheres.
Primitive fitting has also been used for shape abstraction, approximating shapes with cuboids (Zou et al., 2017; Tulsiani et al., 2017) or superquadrics (Paschalidou et al., 2019, 2020) given an input shape. This line of work more commonly take 3D point clouds as input in primitive recovery. Moreover, these works also assume a fixed predefined set of primitives, whereas our work represent shapes as CAD program and thus recovers more general primitives defined by arbitrary closed loop, hence explaining more general objects.
Similar to our CAD representation, learning the primitive decomposition or shape abstractions of a shape allows for intuitive shape editing (Li et al., 2019) and manipulation (Sung et al., 2020). Notably, ReparamCAD (Kodnongbua et al., 2023) is a relevant recent work that learns the variation space given only a single model, though it assumes an underlying shape program (also in the form of boxes) given as input. In contrast, our work learns to output the underlying CAD program given only a single input image without assuming a predefined set of primitives.
@@ 3. Background: Sketch-Extrude CAD
  • $L:(x,y)$, where $(x,y)$ defines the endpoint of a line segment  [^1] Modellers typically assume that the starting point of the first line segment is at the origin..
@@
To result in a valid sketch profile, the resulting curve from the series of commands must be closed and non-self-intersecting  [^2] Without loss of generality, we require sketches to be constructed in a counter-clockwise direction..
Figure 2. An illustrative example of a sketch and extrusion command sequence for a chair’s backrest with a circular cutout. The backrest is created using two extrusion blocks: the first one employs the sketch commands lines and arcs with the extrusion type NewBody, while the second one uses the sketch command circles with the extrusion type Cut.
Extrusion.
@@
The continuous attributes of the extrusion command are given as $E:(\alpha,\theta,\gamma,x,y,z,e)$, where $(\alpha,\theta,\gamma)$ are the three Euler angles defining the extrusion coordinate frame, $(x,y,z)$ denote the origin of the extrusion coordinate system, and $e$ represents the extrusion distance or extent  [^3] Note that in our implementation, there is an additional $b$ argument specifying whether it is a join or cut. As defined, we regard this as two different command types..
@@
4. Method
Reverse engineering a 3D model from input images is a challenging task in computer graphics and vision due to the discrete-continuous nature of the problem. This task requires learning the combinatorial space of shape and program structure along with their corresponding continuous attributes.
@@ 4. Method
Our approach addresses the challenging image-to-CAD problem for common objects by conditionally factorizing the task into two sub-problems. First, we leverage the capabilities of large foundation models (i.e., GPT-4V (Achiam et al., 2023)) to predict the global discrete base structure of the shape from a single input image. This discrete problem involves inferring the semantic parts present in the image and providing the CAD structure, i.e., the command types of each underlying part. Subsequently, we propose a novel transformer-based network called TrAssembler that, conditioned on this semantic discrete base structure, predicts the continuous attributes for the sequence of CAD commands for each semantic part. An overview of the pipeline is shown in Figure 3 and we elaborate on each component in the following subsections.
@@
Figure 3. Overview of our two-stage pipeline. The first stage involves GPT-4V interpreting the input image to decompose it into semantic parts and generate a sequence of discrete CAD structures. The second stage uses a transformer-based model to predict continuous attributes for each part.
4.1. Discrete Global Structure with GPT-4V
In this work, we define the global discrete base structure of a shape as both its part decomposition and the CAD command types for each underlying part. Inferring a shape’s discrete base structure from a single image is challenging due to the combinatorial nature of the problem and the difficulty of generalization across diverse input images from heterogeneous sources. Learning such a model would typically require vast, possibly exponential, amounts of data. Therefore, we leverage large foundation models, specifically GPT-4V (Achiam et al., 2023), which has been trained on extensive amounts of textual and visual data of common objects, such as furniture, to tackle this challenge.
Concretely, we prompt the VLM to infer the discrete global structure by explaining the CAD sketch-extrude language, utilizing in-context learning (Alayrac et al., 2022) by providing a simple example and an additional common object example. Most critically, we employ the chain-of-thought (Wei et al., 2022) technique by asking it to first list the semantic parts from the given image before predicting the CAD sequence for each part.
Furthermore, we prompt the VLM to provide semantic part comments in its predicted global base structure. This involves attaching a part label, i.e., a program comment, before providing the CAD sequence for each identified part. We observe that VLMs’ knowledge of object structure and their ability to perceive it in an image enables consistent labeling regardless of the underlying global structure.
@@ 4.1. Discrete Global Structure with GPT-4V
This semantic labeling is crucial as it facilitates communication and shared learning across corresponding attributes of different objects in our attribute prediction network. This design allows for sharing between shapes with common substructures, even when they have different global discrete base structures.
For example, as shown in Figure 4, backrests from different chairs, recognized and parsed by GPT-4V, are represented within distinct semantic structural blocks, yet the ground-truth CAD commands for these blocks share similar parameters such as axis of extrusion, dimensions, and extrusion distances.
@@
Figure 4. Different chairs still share similar part semantics, and similar parts share similar attribute parameters, facilitating easier attribute prediction.
@@
Figure 5. Overall architecture of TrAssembler. The model first leverages part transformer encoder to generate part embeddings with part structures given by GPT-4V. The part embeddings are then fed into a global transformer encoder and refines them with multi-head self-attention, followed by an MLP to decode the final attribute parameters for each command.

### v2

@@
Abstract.
Reverse engineering 3D computer-aided design (CAD) models from images is an important task for many downstream applications including interactive editing, manufacturing, architecture, robotics, etc. The difficulty of the task lies in vast representational disparities between the CAD output and the image input. CAD models are precise, programmatic constructs that involves sequential operations combining discrete command structure with continuous attributes – making it challenging to learn and optimize in an end-to-end fashion. Concurrently, input images introduce inherent challenges such as photometric variability and sensor noise, complicating the reverse engineering process. In this work, we introduce a novel approach that conditionally factorizes the task into two sub-problems. First, we leverage vision-language foundation models (VLMs), a finetuned Llama3.2, to predict the global discrete base structure with semantic information. Second, we propose TrAssembler that conditioned on the discrete structure with semantics predicts the continuous attribute values. To support the training of our TrAssembler, we further constructed an annotated CAD dataset of common objects from ShapeNet. Putting all together, our approach and data demonstrate significant first steps towards CAD-ifying images in the wild. Code and data can be found in https://github.com/qq456cvb/Img2CAD.
submissionid: 1539journal: TOGccs: Computing methodologies Shape modelingccs: Computing methodologies Shape inferenceccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Secondly, CAD systems, despite their utility in applications requiring detailed editing and manipulation, present significant challenges due to its combination of discrete command types with continuous parameter values. A CAD model is a sequence of discrete geometric operations of different types, e.g. line, arc, circle, extrude add, extrude cut – each characterized by continuous attributes such as length, position, and angle. Learning this dual nature of both discrete operations and their continuous attributes poses a non-trivial challenge, necessitating algorithms to predict both the structural and attribute components accurately. The hybrid discrete-continuous nature of the problem makes image-to-CAD learning non-trivial. This may require an exponential amount of data to learn a combinatorial space of structures, which is difficult to obtain. To tackle these problems, we propose to conditionally factorize the task into two sub-problems – to first predict the global discrete base structure and then, conditioned on this base structure, to regress the continuous attribute values. We leverage the generalizability of large vision foundation models, particularly a finetuned Llama3.2 (Grattafiori et al., 2024), to predict the global base structure from an image. We then propose a transformer-based network named TrAssembler, to regress the continuous attributes conditioned on the discreate structure. While the space of possible object structures is vast, even within a single category such as chairs, shared substructures among objects are common. For example, many chairs have support structures consisting of four legs, and chairs with headrests are more commonly office chairs with five swivel legs. Moreover, we also find that VLM’s can generate semantically consistent part labels as program comments for such common substructures. These labels can then be used by the continuous attribution prediction network to effectively establish correspondences that facilitate shared learning across corresponding attributes (e.g., leg length) of different objects, thus enabling the sharing of information across the attribute space. Our conceptual idea is inspired by the mathematical notion of sheaves (Curry, 2014), where fibers (the attribute parameter spaces in our case) over points in a base space (the space of labelled CAD structures in our case) have partial locally consistent structure (through corresponding attributes). Taking this mathematical concept as inspiration, we leverage shared attributes across different shapes by using shared part names allowing TrAssembler to capture consistency effectively with limited training data.
To further enhance the quality of the generated CAD models, we incorporate recent flow-matching models, specifically GMFlow (Chen et al., 2025b) into our TrAssembler, and employ a masked flow matching loss during training. Given that many human-made objects exhibit strong symmetry priors, we also introduce a symmetry guidance loss during the ODE-based inference sampling. This loss promotes symmetric outputs by performing gradient descent on the estimated score function over the symmetry manifold.
@@
Lastly, to train TrAssembler, we further introduce a CAD dataset of common objects. This was constructed by leveraging the rich part annotations of PartNet (Mo et al., 2019b) and the general assistance of GPT4V (Achiam et al., 2023) – resulting in our CAD-ified subset of ShapeNet (Chang et al., 2015). We will release this dataset upon publication. Our experiments demonstrate the feasibility of our approach for image-to-CAD, which, to the best of our knowledge, is the first of its kind for a general sketch-extrude CAD representation of common objects. We show that our method outperforms baselines and further showcase various applications, such as CAD-ification and shape editing from images in the wild.
@@ 2.1. Reverse Engineering CAD
Reverse engineering has long been a core topic in graphics and CAD (Bénière et al., 2013; Várady et al., 1997; Benko and Varady, 2004; Benkö et al., 2002; Ma et al., 2023). Early methods use inverse CSG modeling (Du et al., 2018; Sharma et al., 2018; Kania et al., 2020; Ren et al., 2021; Yu et al., 2023), but suffer from overly complex decompositions. Recent works leverage large-scale datasets (Koch et al., 2019; Willis et al., 2021; Seff et al., 2020) to enable learning-based methods for B-Rep analysis (Jayaraman et al., 2021; Lambourne et al., 2021; Dupont et al., 2022), surface inference (Guo et al., 2022; Liu et al., 2023), and CAD generation (Khan et al., 2024; Lambourne et al., 2022; Xu et al., 2023; Liu et al., 2024; Chen et al., 2025a; Li et al., 2025). Unlike these mechanical-part-focused datasets, we introduce a dataset of everyday objects with semantic parts.
The sketch-extrude paradigm (Willis et al., 2021; Uy et al., 2022; Xu et al., 2022; Para et al., 2021) supports compact, flexible primitives and has inspired sequential modeling approaches (Wu et al., 2021; Ganin et al., 2021; Jayaraman et al., 2023; Zhang et al., 2024). Point2Cyl (Uy et al., 2022) frames reverse engineering as geometry-aware decomposition, extended by later works to unsupervised (Ren et al., 2022), self-supervised (Li et al., 2023), and multi-view settings (Hong et al., 2024). Our method tackles the harder single-image case.
@@ 2.2. Structured Representation of Shapes
Our daily objects are often highly structured in their geometry and part composition. For example, a chair is composed of parts governed by a set of rules and constraints (Ritchie et al., 2023; Uy et al., 2020, 2021), such as having legs of equal lengths placed symmetrically around the seat. Existing works have explored various ways of representing and learning structure from shapes, such as part-based templates (Ganapathi-Subramanian et al., 2018; Kim et al., 2013; Ovsjanikov et al., 2011), symmetries (Mitra et al., 2006; Sung et al., 2015), shape programs (Müller et al., 2023; Jones et al., 2020, 2021, 2023), and grammars (Kalogerakis et al., 2012; Chaudhuri et al., 2011). Notably, GRASS (Li et al., 2017) and StructureNet (Mo et al., 2019a) represent a shape as a graph of part hierarchies learned through a supervised neural network. The goal of this line of work is to learn higher-level semantic structures, i.e., parts of shapes, but does not focus on the reconstruction of the local geometric detail of each part. These methods typically assume a coarse and homogeneous representation for each part, commonly as bounding boxes. In contrast, our goal is to address image-to-CAD for daily objects, which requires not only understanding the underlying structure but also reconstructing each part’s geometry.
@@ 2.3. Primitive Fitting
Primitive fitting in graphics focuses on decomposing shapes into surface or volumetric primitives. Classical methods detect basic primitives like planes (Czerniawski et al., 2018; Fang et al., 2018; Monszpart et al., 2015) using RANSAC (Schnabel et al., 2007; Li et al., 2011), region growing (Oesau et al., 2016), or Hough voting (Borrmann et al., 2011; Drost and Ilic, 2015). More recent data-driven approaches (Li et al., 2019; Sharma et al., 2020; Lê et al., 2021; Liu et al., 2024) fit primitives such as planes, cylinders, and B-Reps from point clouds using neural networks.
Beyond fitting, primitive-based shape abstraction has also been explored using cuboids (Zou et al., 2017; Tulsiani et al., 2017) or superquadrics (Paschalidou et al., 2019, 2020). A recent work (Jones et al., 2024) also explored using template programs that models shapes as domain-specific programs and represents primitives as cuboids. These methods generally assume 3D point cloud inputs and/or rely on a fixed primitive vocabulary. In contrast, we represent shapes using a CAD program with general primitives defined by arbitrary closed loops, enabling a broader class of objects.
Primitive-based decompositions also support shape editing (Li et al., 2019) and manipulation (Sung et al., 2020). ReparamCAD (Kodnongbua et al., 2023) learns a variation space from a single model but requires a box-based program as input. Our approach instead predicts a CAD program directly from a single image without assuming any pre-defined primitives.
@@ 3. Background: Sketch-Extrude CAD
  • $L:(x,y)$, where $(x,y)$ defines the endpoint of a line segment (modellers typically assume that the starting point of the first line segment is at the origin).
@@
To result in a valid sketch profile, the resulting curve from the series of commands must be closed and non-self-intersecting. Without loss of generality, we require sketches to be constructed in a counter-clockwise direction.
Figure 2. An illustrative example of a sketch and extrusion command sequence for a chair’s backrest with a circular cutout. The backrest is created using two extrusion blocks: the first one employs the sketch commands lines and arcs with the extrusion type NewBody, while the second one uses the sketch command circles with the extrusion type Cut.:
3.2. Extrusion.
@@
The continuous attributes of the extrusion command are given as $E:(\alpha,\theta,\gamma,x,y,z,e)$, where $(\alpha,\theta,\gamma)$ are the three Euler angles defining the extrusion coordinate frame, $(x,y,z)$ denote the origin of the extrusion coordinate system, and $e$ represents the extrusion distance or extent. Note that in our implementation, there is an additional $b$ argument specifying whether it is a join or cut. As defined, we regard this as two different command types.
@@
4. Method: Img2CAD
@@
Reverse engineering a 3D model from input images is a challenging task in computer graphics and vision due to the discrete-continuous nature of the problem. This task requires learning the combinatorial space of shape and program structure along with their corresponding continuous attributes. Our approach addresses this problem for common objects by conditionally factorizing the task into two sub-problems. First, we leverage the capabilities of VLMs (i.e., Llama3.2 (Grattafiori et al., 2024)) to predict the global discrete base structure of the shape from a single input image. This discrete problem involves inferring the semantic parts present in the image and providing the CAD structure, i.e., the command types of each underlying part. Subsequently, we propose a novel transformer-based network called TrAssembler that, conditioned on this semantic discrete base structure, predicts the continuous attributes for the sequence of CAD commands for each semantic part, with flow matching loss and inference time symmetry guidance. An overview of the pipeline is shown in Figure 3 and we elaborate on each component in the following subsections.
@@
Figure 3. Overview of our two-stage pipeline. The first stage involves Llama3.2 decomposing the input image into semantic parts and generate a sequence of discrete CAD structures. The second stage uses a transformer flow matching model to predict continuous attributes for each part.:
4.1. Discrete Global Structure with VLM
In this work, we define the global discrete base structure of a shape as its part decomposition along with the corresponding CAD command types for each constituent part. Inferring a shape’s discrete base structure from a single image is inherently challenging due to the combinatorial complexity and the difficulty of generalizing across diverse input images. To address this, we leverage VLMs, which have been trained on large-scale data. Furthermore, we observe that VLMs such as LLama3.2 or even GPT-4o (Achiam et al., 2023) may sometimes struggle to produce accurate structural predictions, primarily due to their limited inherent understanding of 3D geometry. Hence, we finetune an open-sourced VLM with our training data to achieve better results. Concretely, our VLM uses LLama3.2 finetuned on predicting the discrete base structure of the shapes from our data, i.e. it is not trained on and does not predict the continuous attributes which we leave to our continuous attribute network.
@@ 4.1. Discrete Global Structure with VLM
We prompt the VLM to provide semantic part comments in its predicted global base structure. This involves attaching a part label, before providing the CAD sequence for each identified part. This semantic labeling is crucial as it facilitates the shared learning across corresponding attributes of different objects in our attribute prediction network. This design allows for sharing between shapes with common substructures, even when they have different global discrete base structures. For example, as shown in Figure 4, backrests from different chairs, recognized and parsed by Llama3.2, are represented within distinct semantic structural blocks, yet the ground-truth CAD commands for these blocks share similar parameters such as axis of extrusion, dimensions, and extrusion distances.
@@
Figure 4. Different chairs still share similar part semantics, and similar parts share similar attribute parameters, facilitating easier attribute prediction.:
@@
Figure 5. Overall architecture of TrAssembler. The model first leverages part transformer encoder to generate part embeddings with part structures given by Llama. The part embeddings are then fed into a global transformer encoder and refines them with multi-head self-attention, followed by an MLP to decode the final attribute parameters for each command.: